# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

### v1

@@ 5.3 Time Complexity of Special Cases
In this section, we looked into scheduling for line and out-tree topologies. We illustrate how well the offset schedule approach of Section 3.2 outperforms our proposed methodology in these special cases.
@@
predecessor-flow offset schedule % feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ THEORY 77.8 236.2 0.0 90.1 0 90.1
@@
5.4 Different Criteria
Up until now, we have focused mostly on degeneracy $D_{sum}$ as the criterion. In this section, we briefly study the differences in results when we optimize $D_{sum}({0.75})$ or $D_{\max}$ instead. We ran the local search for instances of set GEN-O.98 with the two aforementioned criteria and compared the results with the baseline results, optimizing $D_{sum}$.
@@ 5.4 Different Criteria
Figure 10: Values of metrics for the instances of GEN-0.98 for different minimization criterion. Instances were sorted in ascending order of $D_{sum}$.
Results per instance are shown in Fig. 10. In each figure, three vertically aligned markers show the values of the metrics for a single instance. Chart (a) shows the values of metrics when $D_{sum}$ was minimized, chart (b) corresponds to the minimization of $D_{sum}({0.75})$, and chart (c) to the minimization of $D_{\max}$.
From all the figures, it is clear that values of $D_{sum}({0.75})$ are highly correlated with the original degeneracy in all experiments, even though the used instances do not necessarily admit solution with $D_{sum}({0.75})=0$ (but they do admit solution with $D_{sum}=0$). The results obtained by minimizing $D_{sum}$ in Fig. 10(a) and $D_{sum}({0.75})$ in Fig. 10(b) are rather similar. However, when $D_{\max}$ is used as an optimization criterion in Fig. 10(c), the worst reported value of $D_{\max}$ among the instances is 8 instead of 9.
Averaged values of the metrics are shown in Table 5. There, the difference between optimizing $D_{sum}$ and $D_{sum}({0.75})$ is clear. Interestingly, direct optimization of $D_{sum}({0.75})$ actually led to an overall worse solution with respect to all the metrics than when we used the $D_{sum}$ criterion, even though the space of the feasible solutions is the same. One of the reasons for this can be the local search procedure; the initial phase, where the chains are reordered, ends after the first failure to improve the criterion. This happens earlier in the case of the $D_{sum}({0.75})$. We believe there might be a connection between the worse performance of optimization with stricter $D_{sum}({\alpha})$ criterion and the difficulty of finding feasible solutions in a TSN setting with release times and deadlines.
criterion metric $D_{sum}$ $D_{sum}({0.75})$ $D_{\max}$ avg $D_{sum}$ 143.4 175.6 206.5 avg $D_{sum}({0.75})$ 274.9 296.7 347.8 avg $D_{\max}$ 5.07 5.35 4.70
Table 5: Average values of metrics given the optimization criterion. Results reported for instance set GEN-0.98.
5.5 Real Life-like Instances
Finally, we performed several experiments on real-life inspired instances, which highlight how well our approach performs in setting much closer to the practical applications.
@@ 5.5.1 Production Line
Firstly, we used production-line instances from [29]. In that paper, three types of instances were generated, one of which used harmonic periods. The original paper handled problems with more complex instances than we do, including frame-isolation constraints, queues, release times, and deadlines, and reported low success rates once the maximum utilization among the resources rose above 50%. These additional constraints are omitted for the purpose of this paper.
% feasible avg $D_{sum}$ 0-$D_{sum}$ production-line 100.0 530.9 80.5
@@
In Table 6, we present the results across the 3000 mentioned instances, obtained after six minutes. For most instances, we find a 0-$D_{sum}$ solution. For the rest of the instances, the values of degeneracy were quite high. However, in these cases, the number of tasks was, on average, 30000, and the number of chains was 2700. The average $D_{\max}$ was equal to $6.2$, i.e., the most delayed chain reported degeneracy $<7$. Altogether, better results could be achieved by providing the solver more time, as in these cases, optimization has not yet converged. However, for these large-scale instances, it would probably be worth investigating more aggressive neighborhood operators or decomposition techniques. Note that CP warm start was still able to determine feasible schedules even for these large instances.
@@ 5.5.2 Avionic
The second real-life set is the avionic instances, also from [29]. Chains of the five industrial instances from the avionic domain were scheduled across 70-92 resources, with a maximum utilization of less than 38%. Thus, we also created a second set of instances in which the periods were halved, doubling the utilization. However, in both cases, our approach found a feasible 0-$D_{sum}$ solution within two minutes (given the constraints we study in this paper).
@@ 5.5.3 Automotive
Finally, we generated a set of automotive instances following a zonal architecture in modern cars. An example of such architecture is shown in Fig. 11(a), which we used as an inspiration for our network in Fig. 11(b). The central controller used 5Gb/s communication speed, the encompassing ring used 2.5Gb/s, and the rest of the connections used 1 Gb/s. We generated messages between leaf and central nodes of a size from 72 to 1542 bytes, as is used in Ethernet frames. The messages’ periods were 10 ms, 50 ms, 250 ms, and 500 ms.
@@
We generated 20 such instances with an average maximum utilization of $U_{m}=0.9$ and 300000+ tasks each. However, due to the maximum length of the Ethernet messages, their processing times were short in contrast to the minimum period. Thus, our approach found feasible solutions with $D_{\max}\leq 2$ within 6 minutes, even for these large-scale instances. Even though it managed to perform only a few iterations of the local search, it managed to produce a schedule with satisfactory characteristics.
@@ 6 Conclusion
In this paper, we presented complexity results related to the periodic scheduling problem with harmonic periods and dedicated resources. Crucially, we have shown that for communication over the line and tree topologies with a single source node, an offset schedule approach can be used to efficiently find a high-quality solution.
We proposed a warm start approach, which successfully helps the local search optimize instances with very high utilization. Furthermore, we described a CP matheuristic, which leverages CP scheduling capabilities by optimizing the schedule while targeting one period at a time.
Several interesting conclusions can be drawn from our results. Even though scheduling periodic tasks is the principal component of TSN, we achieved much higher utilization rates in our restricted problem than is typical in TSN applications; it seems that the additional constraints introduced by the TSN standard are the main culprit behind the difficulty of scheduling, not just the PSP alone.
Another interesting property is the relationship between the number of tasks and the difficulty of solving the instance. We had problems even generating instances that were not quickly solved by the baseline heuristic. For a fixed size of the ”observational interval”, increasing the number of tasks will, at one point, necessarily decrease the average processing time of tasks, which generally makes them easier to schedule.
Finally, we successfully demonstrated how understanding the instances’ characteristics can be exploited in the PSP domain. To name a few, using specialized methods for specific resource topologies or focusing on optimizing an isolated sub-problem, which was, in our case, a set of tasks corresponding to a single period.
@@ Acknowledgments
This work was supported by the Grant Agency of the Czech Republic under the project GACR 22-31670S and by the Grant Agency of the Czech Technical University in Prague, grant No. SGS22/167/OHK3/3T/13.

### v2

@@ 5.3 Time Complexity of Special Cases
In this section, we looked into scheduling for acyclic instances with a bottleneck resource, specifically instances with line and out-tree support graphs. We illustrate how well the offset schedule approach of Section 3.2 outperforms our proposed methodology in these special cases.
@@
predecessor-flow offset schedule % feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ feasible $D_{sum}$ $D_{sum}^{chains}$ 0-$D_{sum}$ THEORY 77.8 114.0 (335.2) 2.8 (2.5) 0.0 90.1 0.0 (0.0) 0.0 (0.0) 90.1
@@
5.4 Different Objectives
Up until now, we have focused mostly on degeneracy $D_{sum}$ as the objective. In this section, we briefly study the differences in results when we optimize $D_{sum}({0.75})$ or $D_{\max}$ instead. We ran the local search for instances of set GEN-0.98 with the two aforementioned objectives and compared the results with the baseline results, optimizing $D_{sum}$.
objective metric $D_{sum}$ $D_{sum}({0.75})$ $D_{\max}$ $D_{sum}$ 90.0 (181.1) 110.0 (228.5) 124.5 (259.2) $D_{sum}(0.75)$ 191.0 (280.3) 191.5 (340.6) 251.0 (357.3) $D_{\max}$ 5.0 (1.7) 5.0 (1.4) 5.0 (1.5)
Table 5: Average values of metrics given the objective. Results reported for set GEN-0.98.
@@ 5.4 Different Objectives
Figure 11: Values of metrics for the instances of GEN-0.98 for different objectives. Instances were sorted in ascending order of $D_{sum}$.
Results per instance are shown in Fig. 11. In each figure, three vertically aligned markers show the values of the metrics for a single instance. Fig. 11(a) shows the values of metrics when $D_{sum}$ was minimized, Fig. 11(b) corresponds to the minimization of $D_{sum}({0.75})$, and Fig. 11(c) to the minimization of $D_{\max}$.
From all the figures, it is clear that values of $D_{sum}({0.75})$ are highly correlated with the original degeneracy in all experiments, even though the used instances do not necessarily admit solution with $D_{sum}({0.75})=0$ (but they do admit solution with $D_{sum}=0$). The results obtained by minimizing $D_{sum}$ in Fig. 11(a) and $D_{sum}({0.75})$ in Fig. 11(b) are rather similar. However, when $D_{\max}$ is used as an objective in Fig. 11(c), the worst reported value of $D_{\max}$ among the instances is 8 instead of 9.
Median and standard deviations of the metrics are shown in Table 5. There, the difference between optimizing $D_{sum}$ and $D_{sum}({0.75})$ is clear. Interestingly, direct optimization of $D_{sum}({0.75})$ led to an overall slightly worse solution concerning all the metrics than when we used the $D_{sum}$ objective, even though the space of the feasible solutions is the same. One reason for this is the local search procedure; the initial phase, where the chains are reordered, ends after the first failure to improve the objective function. This happens earlier in the case of the $D_{sum}({0.75})$. We believe there might be a connection between the worse performance with stricter $D_{sum}({\alpha})$ objective and the difficulty of finding feasible solutions in a TSN setting with release times and deadlines.
5.5 Real-Life Inspired Instances
Finally, we performed several experiments on real-life inspired instances.
@@ 5.5.1 Production Line
Firstly, we used production-line instances from [4]. In that paper, three types of instances were generated, one of which used harmonic periods. The original paper handled problems with more complex instances than we do, including frame-isolation constraints, queues, release times, and deadlines, and reported low success rates once the maximum utilization among the resources rose above 50%. These additional constraints are omitted in this paper.
% feasible avg $D_{sum}$ avg $D_{sum}^{chains}$ 0-$D_{sum}$ production-line 100.0 530.9 $\pm$ 1332.5 0.1 $\pm$ 0.1 80.5
@@
In Table 6, we present the results across the 3000 mentioned instances, obtained after six minutes. We report average values and standard deviations of the degeneracy metrics since their medians were 0. For most instances, we find a 0-$D_{sum}$ solution. For the rest of the instances, the degeneracy was high. However, in these cases, the number of tasks was, on average, 30000, and the number of chains was 2700. The average $D_{\max}$ was equal to $6.2$, i.e., the most delayed chain reported degeneracy $<7$. Altogether, better results could be achieved by providing the solver more time, as optimization has not yet converged in these cases. However, it would probably be worth investigating more aggressive neighborhood operators or decomposition techniques for these large-scale instances. Note that CP warm start could still determine feasible schedules even for these large instances.
@@ 5.5.2 Avionic
The second real-life set is the avionic instances, also from [4]. Chains of the five industrial instances from the avionic domain were scheduled across 70-92 resources, with a maximum utilization of less than 38%. Thus, we also created a second set of instances in which the periods were halved, doubling the utilization. However, for all 10 instances, our approach found feasible 0-$D_{sum}$ solutions within two minutes (given the constraints we study in this paper).
@@ 5.5.3 Automotive
Finally, we generated a set of automotive instances following a zonal architecture in modern cars. An example of such architecture is shown in Fig. 12(a), which we used as an inspiration for our network in Fig. 12(b). The central controller used 5Gb/s communication speed, the encompassing ring used 2.5Gb/s, and the rest of the connections used 1 Gb/s. We generated messages between leaf and central nodes with between 72 to 1542 bytes, as is used in Ethernet frames, and determined the processing times of the jobs based on both the message size and the resource’s speed. The messages’ periods were 10 ms, 50 ms, 250 ms, and 500 ms.
@@
We generated 20 such instances with an average maximum utilization of $U_{m}=0.9$ and 300000+ tasks each. However, due to the maximum length of the Ethernet messages, their processing times were short in contrast to the minimum period. Thus, our approach found feasible solutions with $D_{\max}\leq 2$ within 6 minutes, even for these large-scale instances. Even with just a few iterations, the local search produced a schedule with satisfactory characteristics.
@@ 6 Conclusion
In this paper, we presented complexity results related to the periodic scheduling problem with harmonic periods and dedicated resources. Crucially, we have shown that for acyclic cases with a bottleneck resource, the offset schedule approach can be used to find a high-quality solution efficiently.
We proposed a warm start approach, which successfully helps the local search optimize instances with high utilization. Furthermore, we described a CP matheuristic, which leverages CP scheduling capabilities by optimizing the schedule while targeting one period at a time.
Several interesting conclusions can be drawn from our results. Even though scheduling periodic tasks is the principal component of TSN, we achieved much higher utilization rates in our restricted problem than is typical in TSN applications; it seems that the additional constraints introduced by the TSN standard are the main culprit behind the difficulty of scheduling.
Another interesting property is the relationship between the number of tasks and the difficulty of solving the instance. We had problems generating instances that were not quickly solved by the baseline heuristic. For a fixed size of the “observational interval”, increasing the number of tasks will, at one point, necessarily decrease the average processing time of tasks, which generally makes them easier to schedule.
Finally, we successfully demonstrated how understanding the instances’ characteristics can be exploited in this domain. To name a few, using specialized methods for specific underlying support graphs or focusing on optimizing an isolated sub-problem, which was, in our case, a set of tasks corresponding to a single period.
@@ Acknowledgments
This work was supported by the Grant Agency of the Czech Republic under the project GACR 22-31670S and under the project GACR 25-17904S.
@@
List of Main Symbols
Symbol Description Symbol Description
[…]
$\tau(\beta)$ $\beta$-th harmonic period $\hat{H}$ hyper-period
$w$ least period $T^{k}$ period of $\mathcal{C}^{k}$
${C}^{\tau(\beta)}$ set of chains with period $\tau(\beta)$ $J_{i}^{k}$ task $i$ of $\mathcal{C}^{k}$
$n^{k}$ number of tasks in $\mathcal{C}^{k}$ $p_{i}^{k}$ processing time of $J_{i}^{k}$
$\pi_{i}^{k}$ assignment of $J_{i}^{k}$ to resource $t_{i}^{k}$ start time of first occurrence of $J_{i}^{k}$
$\sigma_{i}^{k}$ start time of $J_{i}^{k}$ in core schedule $s_{i}^{k}$ start time of $J_{i}^{k}$ in shifted core schedule
$o_{m}$ offset of resource $m$ $U_{m}$ utilization of resource $m$
$U$ utilization of the instance $H$ height of bin
$R_{i}^{k}$ rectangle associated with $J_{i}^{k}$ $\ell_{i}^{k}$ width of $R_{i}^{k}$
$0pt{i}{k}$ height of $R_{i}^{k}$ $J$ ordered list of tasks
$S^{k}$ start-to-end latency of $\mathcal{C}^{k}$ $D^{k}$ degeneracy of $\mathcal{C}^{k}$
$D^{k}(\alpha)$ $\alpha$ degeneracy of $\mathcal{C}^{k}$ $D_{sum}$ degeneracy objective
$D_{\max}$ maximum degeneracy objective $D_{sum}({\alpha})$ $\alpha$ degeneracy objective